OpenStudy (anonymous):

5x-2 divided by 7 = x+3 divided by 2

OpenStudy (anonymous):

what does it equal

OpenStudy (anonymous):

Are you finding x? cross mutiply it

OpenStudy (anonymous):

yes X

OpenStudy (anonymous):

if a/b = c/d, then ad=bc use this property, u wil get a nice linear function to solve.